OpenStudy (anonymous):

given log6(4)=X, log6(10)=y, log6(9)=z then find log6(1/25) in terms of X, Y & Z

OpenStudy (anonymous):

is the answer -2Y+\[\sqrt{?}\]

OpenStudy (anonymous):

I mean -2y + squareroot(x)?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

I don't know. I'm trying to figure this out.

OpenStudy (anonymous):

ok here is how i did it: I am not writing base since it is 6. I will just write numbers involved.. log(1/25) = log(5^-2) = -2*log5 ----- eqn (i) [so log5 is what you are looking for] log10 = log(2*5) = log2 + log5 = y log5 = y - log2 ---- eqn (ii) .. [so you will need value of log2 to solve log5] log4 = log(2^2) = 2*log2 = x log2 = x/2 --- equation(iii) finally combining all back to equation i log(1/25) = -2*log5 = -2 * ( y - log2) [from equation ii put value for log 5) =-2 * (y-x/2) -- [log2 value comes form equation 3) = -2y + x .. this is your final answer .. I screwed up when i said the answer earlier ..

OpenStudy (anonymous):

Wow. That's great. I'll try to work through it to make sure I understand it. But I think I got it. Thanks! U da man. Unless you're a woman. Either way, thanks again!

OpenStudy (anonymous):

hahah .. I am a man buddy!! .. thanks for da medal ...